OpenStudy (anonymous):

Daniel and Manuel have $600 altogether. If Daniel spends 2/5 of his money and Manuel spends $40, then they will have the same amount of money left. how much money does Daniel have?

OpenStudy (anonymous):

2/5x + 40 = 600 -40 -40 2/5x = 560 2/5 is 0.4 0.4x = 560 /0.4 /0.4 x=1400 Daniel had $1,400.

OpenStudy (anonymous):

Comprende? :)

OpenStudy (anonymous):

um i respect your help sathelp but it says in my problem that Daniel and Manuel have $600 together, so $1400 is the wrong answer :/

OpenStudy (anonymous):

No, Daniel spent 2/5 of $1,400 which equals $560 plus Manuel's money in which he spent $40 which would give us $600. He had 1,400 originally but 2/5 of that would be $560. So, your right, it would be $560. Thanks for the heads up.

OpenStudy (anonymous):

I think its asking for Daniel's money in total though so I think it is $1,400 :/